a seller claim that her lip tint has a mean organic content of 90%. A rival seller asked 60 users of that lip tint and found that it has a mean organic content of 85% with a standard deviation of 5%. test the claim at 1% level of significance and assume tha the population is approximately normally distributed



1
Expert's answer
2022-05-19T18:50:21-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=90"

"H_a:\mu\not=90"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=59" degrees of freedom, and the critical value for a two-tailed test is "t_c =2.661759."The rejection region for this two-tailed test is "R = \{t:|t|> 2.661759\}."

The t-statistic is computed as follows:



"t=\dfrac{85-90}{5/\sqrt{60}}=-7.746"

Since it is observed that "|t| = 7.746>2.661759=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed "df=59" degrees of freedom, "t=-7.746" is "p=0," and since "p=0<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than 90, at the "\alpha = 0.01" significance level.